if m= -2/7 and is the slope thru (-8,r) and (4,3) what is r?
Answer by lwsshak3(11628)

if m= -2/7 and is the slope thru (-8,r) and (4,3) what is r?
**
slope=∆y/∆x=(y1-y2)/(x1-x2)
=(3-r)/(4-(-8))
=(3-r)/(4+8)
=(3-r)/12=-2/7
21-7r=-24
7r=45
r=45/7